OpenStudy (anonymous):

For f(x)=1/x-5 and g(x) = x^2+2, find (fog)(x), and (gof)(6)

OpenStudy (anonymous):

fog=f(g(x)) so it pretty much means put g(x) where ever x appears in f(x) so fog=1/(x^2+2)-5

OpenStudy (anonymous):

so first one is fog=1/(x^2+2)-5, what about 2?

OpenStudy (anonymous):

do the same thing g(f(x))=gof so everywhere theres and in in g(x) put f(x). And I'll let you do that